OpenStudy (anonymous):

On the graph of the equation 2x − y = 10, what is the value of the y-intercept?

OpenStudy (bradely):

2x-y=10 for y intercept x=0 2(0)-y=10 -y=10 y=-10 Source: http://www.mathskey.com/question2answer/